Just as her parachute opens, a 60-kg parachutist is falling at a speed of 50.0 m/s. After 0.80 seconds has passed, the chute is fully open and her speed has dropped to 12.0 m/s. Find the average retarding force exerted upon the chutist during this time if the deceleration is uniform.

1
Expert's answer
2021-12-02T10:03:45-0500

"v=v_0-at\to a=\frac{v_0-v}{t}=\frac{50-12}{0.8}=47.5\ (m/s^2)"


"F=ma=60\cdot47.5=2850\ (N)" . Answer